A ball is dropped from a height h above the ground. Neglect the air resistance, it's velocity v varies with its height y above the ground as
a) √2g(h-y)
b) √2gh
c)√2gy
d)√2g(h+y)

Grade:12th pass

1 Answers

Harshit Goyal
6 years ago
if height from which it has been dropped is `h` and for a point `y` m above the ground, the distance travelled by the ball will be (h-y). Answer - option AV² - U² = 2 * (-g)*(-(h-y))Since it has been dropped U = 0 and simplifying we get, V² = 2g(h-y)Hence V = √(2g(h-y))